# XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge

### v2

@@ 4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs under cross-modality FSL settings. Three stages are designed, as shown in Figure 1. In the Splice stage, default reshaping (see Appendix B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features, especially when layer shapes vary during layer-wise search. We propose a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structural gaps. In the Repair stage, each connector is fine-tuned by the generative transfer module to generate adaptive channels, preserve key sensing features, and minimize MMC shifts. Finally, the Removal stage applies channel removal (see Appendix B.2) to further enhance layer-wise performance.
@@ 4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability, our intuition is to safely fine-tune few sensor data by aligning the misinterpreted latent features with the original features (i.e., anchors) to minimize MMC shifts. We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant or noisy ones.
Reduced-orthogonal feature space. We project anchor MMCs into a lower-dimensional, orthogonal feature space using principal component analysis (PCA) (Valpola, 2015) (i.e., anchor PCA space). Unlike the raw MMC feature space, where channel dimensions are often noisy and correlated, PCA enforces orthogonality across components, removing redundancy and suppressing noise. This enables a compact subspace (i.e., defaulting to 2 components) where the projected data captures maximum variance and forms highly clustered distributions. The optimized component weights (i.e., projection coefficients) highlight the most important channels contributing to layer performance. Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 1(a). Once initialized for each layer, this anchor PCA space is reused to project sensing MMCs.
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 1(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen & Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
$\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
@@
Anchor class pairing. Building on the key finding that class pairing boosts latent feature alignment across modalities (see Section 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the anchor-based alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
@@ 4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the anchor-based alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Figure 1). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the distance ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
@@ 5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) effectively select and recombine layers of interest to form a compact model.
@@ 5.1 Search control design
Following NAS principles (White et al., 2023), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To effectively manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the highest-valued repaired layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer to maintain layer-wise dependence. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is device-specific resource budget.
@@ 5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see Appendix B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in Appendix B.5).
6 Experiment
@@
  TL (Jiang et al., 2022) Single-source A standard transfer learning method by freezing all layers
FT (Oh et al., 2022) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method
MAML (Chen et al., 2019) Single-source A learning-based FSL method
DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method using adversarial ProtoNet
SemiCMT (Chen et al., 2024) Single-source A cross-modal method using contrastive alignment
GPT2 (Zhou et al., 2023) Single-(LLM) A large language model based method using fine-tuning
MDDA (Zhao et al., 2020) Multi-source A cross-domain FSL method using knowledge distillation
MCW (Lee et al., 2019) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
@@
  Dataset Modality Subject Class Type Shape Privacy
@@
miniImageNet (Vinyals et al., 2016) Image - 100 - 3x84x84 Public
miniDomainNet (Zhou et al., 2021) Image - 126 - 3x224x224 Public
Office-31 (Saenko et al., 2010) Image - 31 - 3x224x224 Public
Office-Home (Venkateswara et al., 2017) Image - 65 - 3x224x224 Public
Caltech-101 (Fei-Fei et al., 2004) Image - 101 - 3x224x224 Public
Newsgroup (Thulasidasan et al., 2021) Text - 20 - 100x1x1000 Public
VoxCeleb (Nagrani et al., 2017) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen et al., 2010) IMU 12 11 Time-S 77x1x100 Public
@@
HHAR (Stisen et al., 2015) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt et al., 2018) ECG 15 3 Time-S 10x1x256 Public
Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
Writing Ultrasound 10 10 Time-S 1x1x256 Private
Emotion mmWave 10 7 Spectrum 3x224x224 Private
BP mmWave 10 - Doppler 1x1x499 Private
ChestX (Wang et al., 2017) Image - 5 - 1x1024x1024 Public
@@
Table 3: DL model backbone specifications.
6.1 Experiment setup
We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1). Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3). To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi). Based on these, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (see Appendix A for details).

### v3

@@ 4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, restoring pre-trained layer performance under cross-modality FSL settings. As shown in Figure 3, SRR comprises three stages. Splice: instead of default reshaping using fixed up/down-sampling (Appendix B.1), which is non-trainable and can discard features, especially when layer shapes vary during layer-wise search, we propose a trainable, compact connector placed between heterogeneous layers to enforce shape compatibility, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair. Repair: each connector is fine-tuned by the generative transfer module to adapt layer channels and minimize MMC shifts. Removal: layer channel removal (Appendix B.2) is applied to further strengthen repairing.
@@ 4.1 Repairing in complex feature space
Since source layers are already highly discriminative, our intuition is to safely adapt few sensor data by aligning their misaligned latent feature distributions to the "original" source-layer distributions, which serve as anchors. We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. Since the raw MMC feature space is high-dimensional, noisy and correlated across channels, our insight is to reduce MMC dimensionality by preserving key layer channels while suppressing noise.
Reduced-orthogonal feature space. For each layer, we project anchor MMCs into a low-dimensional, orthogonal feature space via PCA (Valpola, 2015) (i.e., anchor PCA space). PCA removes redundancy by decorrelating channel dimensions and suppresses noise, yielding a compact subspace (we use 2 components by default) that preserves maximal variance and typically forms highly clustered class distributions. The resulting projection coefficients (i.e., component weights) also highlight the most important layer channels that contribute to layer performance. Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 3(a). Once initialized per layer, the same anchor PCA space is reused to project sensing MMCs.
Feature space alignment. After projection, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 3(a). Due to MMC shifts, the scale and orientation of Sensing-center are notably different from those of Anchor-center. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen and Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
$\scriptsize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
@@
Anchor class pairing. Building on the key finding that class pairing boosts latent feature alignment across modalities in Figure 2(b), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes by minimizing a pairing shift objective (i.e., sum of centroid distances across pairs). This is solved as a linear sum assignment via the standard Hungarian algorithm, outputting the optimal pairing set $pairST$. This pairing is fully optimized by similarity in the given feature space.
@@ 4.2 Repairing layer-wise dependence
Generative transfer module. To restore layer-wise dependence, we design a generative transfer module that optimizes the generator (i.e., connector) using an adversarial objective. The discriminator operates on each frozen pre-trained layer together with its anchor PCA space (Figure 3). Freezing preserves stable anchors and prevents overfitting. Given each pairing set $pairST$, the connector is fine-tuned to minimize MMC shifts, uniquely transforming sensor inputs to align with anchors.
Anchor-based repair loss. We propose an anchor-based repair loss ($loss_{srr}$) to harmonize layer-wise latent feature distributions across modalities in the anchor PCA space. For each class pair $(c_{s},c_{t})\in pairST$, our objective (i) minimizes the distance $D(\cdot)$ between the centroids of the projected anchor and sensing MMC distributions (i.e., positive loss), and (ii) maximizes the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. The loss function is formulated as:
$\scriptsize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
@@ 5 Layer Recombining
Working with the SRR pipeline, XTransfer performs layer recombining via LWS control to (i) perform the pre-search check that accelerates the search, (ii) stabilize the search using the dynamic search range mechanism, and (iii) select and recombine layers of interest to restructure models.
@@ 5.1 Search control design
Inspired by NAS (White et al., 2023), effective layer-wise search relies on properly designing the search space, actions, and strategy under on-device resource constraints. We hence define the search space as all layers from source models, organized into search pools containing layer candidates (i.e., repaired layers). Each pool forms a window of size $I\times J$, where $I$ is the number of source models (width) and $J$ is the number of L-units (depth). We also define four search actions: init selects a starting layer (default depth $1$); continue and skip perform intra-model recombination; and cross enables inter-model recombination.
Resource-constrained search strategy. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ is the connector and $L_{ij}$ is each repaired layer. Under resource constraints, LWS control selects the highest-valued candidate in each pool and verifies that it improves layer-wise dependence by checking whether its S-score exceeds that of the previously selected layer. If selected, $L_{ij}$ becomes a layer of interest and is recombined, and the next pool starts from layer index $j+1$. Otherwise, the pool is discarded, and the window shifts forward. To incorporate the constraints, we define a resource-constrained function formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$ where $n$ is the current recombined layer index and $L$ is the maximum number of L-units. Each candidate’s adjusted overhead is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ measures actual cost (e.g., weighted FLOPs and memory), and $Max(Res(P))$ is the maximum cost in current pool $P$. The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is device-specific resource budget.
@@ 5.2 Efficient search
Pre-search check. Although SRR fine-tunes connectors efficiently, applying SRR to every candidate in a large pool can still slow down LWS control. We hence propose the pre-search check strategy that avoids unnecessary repairs before applying SRR. We track S-score progression across three states: after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After). Empirically, SRR improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation (see Appendix B.3 for details). Based on this insight, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (Appendix B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. Estimation starts after init and is updated after each newly selected layer.
Dynamic search range mechanism. Using $rate_{n}^{est}$, LWS can initially apply a narrow search range (e.g., only the top-1 candidate by estimated repair rate) to filter most candidates before SRR. However, since $rate_{n}^{est}$ is updated sparsely, estimation errors may cause missed optimal candidates. To stabilize search while retaining efficiency, we develop a dynamic search range mechanism that adjusts the search range, which is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range. Here, $\hat{u}$ is a signed unit vector. If $rate_{n-1}^{est}<rate_{n-1}$, it indicates that the estimation might be inaccurate, and $\hat{u}$ will be positive to increase the search range to repair more candidates and stabilize accuracy. Otherwise, the range will be decreased to bypass unnecessary repairs and improve efficiency.
@@
  Transfer Learning (TL) (Jiang et al., 2022) Single-source A standard transfer learning method that reuses a source model by freezing all layers.
Fine-tuning (FT) (Oh et al., 2022) Single-source A conventional fine-tuning method that updates all layers of a pre-trained model using available target data.
ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method that constructs class prototypes by averaging support set and classifies query set in the feature space.
MAML (Chen et al., 2019) Single-source A gradient-based meta-learning method that learns a source model with shared initialization parameters from scratch.
DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method that uses adversarial ProtoNet with a domain discriminator to reduce latent feature distribution shift.
SemiCMT (Chen et al., 2024) Single-source A cross-modal learning method that applies contrastive self-supervised learning during model adaptation.
GPT2 (Zhou et al., 2023) Single-(LLM) A LLM-based method that enables in-context learning after fine-tuning on few target data.
MDDA (Zhao et al., 2020) Multi-source A multi-domain FSL method that uses adversarial discriminative adaptation and multi-source knowledge distillation.
MCW (Lee et al., 2019) Multi-source A multi-domain FSL method that uses maximal correlation analysis to fuse multiple source domains for few-shot model adaptation.
MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML that uses extensive target data to train from scratch, serving as an oracle upper bound.
6 Experiment
6.1 Experiment setup
@@
  Dataset Modality Subject Class Train Size Type Shape Privacy
@@
miniImageNet (Vinyals et al., 2016) Image - 100 50k - 3x84x84 Public
miniDomainNet (Zhou et al., 2021) Image - 126 137.5k - 3x224x224 Public
Office-31 (Saenko et al., 2010) Image - 31 3.3k - 3x224x224 Public
Office-Home (Venkateswara et al., 2017) Image - 65 12.7k - 3x224x224 Public
Caltech-101 (Fei-Fei et al., 2004) Image - 101 7.3k - 3x224x224 Public
Newsgroup (Thulasidasan et al., 2021) Text - 20 15k - 100x1x1000 Public
VoxCeleb (Nagrani et al., 2017) Audio - 100 14.9k Time-S 1x512x300 Public
OPPORTUNITY (Roggen et al., 2010) IMU 12 11 3.7k Time-S 77x1x100 Public
@@
HHAR (Stisen et al., 2015) IMU 9 6 - Time-S 6x1x256 Public
WESAD (Schmidt et al., 2018) ECG 15 3 - Time-S 10x1x256 Public
Gesture IMU/PPG 10 8 - Time-S 7x1x200 Private
Writing Ultrasound 10 10 - Time-S 1x1x256 Private
Emotion mmWave 10 7 - Spectrum 3x224x224 Private
BP mmWave 10 - - Doppler 1x1x499 Private
ChestX (Wang et al., 2017) Image - 5 - - 1x1024x1024 Public
We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1). Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3). To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi, Appendix Table 5 for details).
Based on the testbeds, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (Appendix A for more details on our experiment setup).
Table 3: DL model backbone specifications.